# Double-coset zeta functions for groups acting on trees

### v1

@@ 2.1. Generalities
Whenever there is no ambiguity, we denote a $1$-point set $\{x\}$ by the element $x$ itself. Given a set $X$, for every subset $A\subseteq X$ denote by $\mathbbm{1}_{A}$ the indicator function of $A$. For every set $X$, the symmetric group $\mathrm{Sym}(X)$ is always regarded as a topological group with the permutation topology, i.e., the topology generated by the local basis at $1$ given by all possible pointwise stabilisers of finite subsets of $X$.
@@ 2.2. Graphs
A graph (in the sense of J-P. Serre [20]) consists of a set $\Gamma=V\Gamma\sqcup E\Gamma$ partitioned into two subsets $V\Gamma$ and $E\Gamma$ (called the set of vertices and the set of edges of $\Gamma$, respectively), together with two maps $o,t\colon E\Gamma\to V\Gamma$ (called origin and terminus maps, respectively) and an involution ¯$\overline{\cdot}\colon E\Gamma\to E\Gamma$ (called edge inversion) satisfying $\bar{e}\neq e$ and $o(\bar{e})=t(e)$, for every $e\in E\Gamma$. We introduce the following notation for a graph $\Gamma$:
@@
An orientation in a graph $\Gamma$ is a set $E\Gamma^{+}\subseteq E\Gamma$ satisfying |$|\{e,\bar{e}\}\cap E\Gamma^{+}|=1$ for every $e\in E\Gamma$, and $E\Gamma=E\Gamma^{+}\cup\{\bar{e}\mid e\in E\Gamma^{+}\}$. A graph $\Gamma$ is non-empty if it has at least one vertex, and is locally finite if $|o^{-1}(c)|<\infty$ for all $c\in V\Gamma$. A subgraph of a graph $\Gamma$ is a subset $\Lambda$ such that $o(E\Gamma\cap\Lambda),t(E\Gamma\cap\Lambda)\subseteq V\Gamma\cap\Lambda$ and $\overline{e}\in E\Gamma\cap\Lambda$ for every $e\in E\Gamma\cap\Lambda$. The subset $\Lambda$ inherits a graph structure from $\Gamma$. A subgraph $\Lambda$ of $\Gamma$ is proper if $\Lambda\neq\Gamma$. A vertex $v$ in $\Gamma$ is said to be terminal if $|o^{-1}(v)|=1$. An edge $e$ in $\Gamma$ with $o(e)=t(e)$ is called $1$-loop. A $n$-bouquet of loops (based at $c$) is a graph with one vertex $c$ and edge-set $\{a_{i},\bar{a}_{i}\mid 1\leq i\leq n\}$, where each $a_{i}$ is a $1$-loop starting at $c$. A $1$-segment is a graph $\Gamma$ with two distinct vertices and an edge-couple $\{e,\bar{e}\}$ connecting them.
@@
Let $\Gamma$ be a graph. A path in $\Gamma$ is a sequence of vertices and edges $\mathfrak{p}=(v_{0},e_{1},v_{1},\ldots,e_{n},v_{n})$, $n\geq 0$, with $o(e_{i})=v_{i-1}$ and $t(e_{i})=v_{i}$ for every $1\leq i\leq n$. We say that $\mathfrak{p}$ starts at $v_{0}$ (or at $e_{1}$) and ends at $v_{n}$ (or at $e_{n}$), has reverse path is $\overline{\mathfrak{p}}=(v_{n},\bar{e}_{n},v_{n-1},\ldots,\bar{e}_{1},v_{0})$, and length $n$ (written $\ell(\mathfrak{p})=n$). If $n\geq 1$, we may without ambiguity specify only the sequence of edges. If $n=0$, the $1$-term sequence $\mathcal{O}_{v_{0}}=(v_{0})$ is called the trivial path at $v_{0}$. Denote by $\mathcal{P}_{\Gamma}$ the set of all paths in $\Gamma$. Given non-empty subsets $X,Y\subseteq\Gamma$, let $\mathcal{P}_{\Gamma}(X\to Y)$ be the set of all paths in $\Gamma$ starting at some $x\in X$ and ending at some $y\in Y$. The product of two paths $\mathfrak{p}=(e_{1},\ldots,e_{m})$ and $\mathfrak{q}=(f_{1},\ldots,f_{n})$ is defined only if $t(e_{m})=o(f_{1})$ and it is the path $\mathfrak{p}\cdot\mathfrak{q}=(e_{1},\ldots,e_{m},f_{1},\ldots,f_{n})$. If $\mathfrak{p}$ is a path starting and ending at the same vertex, denote by $\mathfrak{p}^{d}$ the $d$-th power of $\mathfrak{p}$ with respect to the product defined before. A path $\mathfrak{p}$ is reduced if either $\ell(\mathfrak{p})=0$ or $\mathfrak{p}=(e_{1},\ldots,e_{n})$ and $e_{i+1}\neq\bar{e}_{i}$ for every $11\leq i\leq n-$ 1. For $n\geq 1$, an $n$-cycle is a reduced path $\mathfrak{p}=(e_{1},\ldots,e_{n})$ with $o(e_{1})=t(e_{n})$ and $t(e_{i})\neq t(e_{j})$ for all $11\leq i,j\leq n-$ 1 with $i\neq j$.
@@
A graph $\Gamma$ is connected if for all $v,w\in V\Gamma$ there is a path from $v$ to $w$. A subgraph of $\Gamma$ is a connected component if it is a maximal connected subgraph of $\Gamma$. A graph is the disjoint union of all its connected components. A tree $T$ is a connected graph with no $n$-cycles, for every $n\geq 1$. If $T$ is a tree and $e\in ET$, then the graph $T\setminus\{e,\bar{e}\}$ has two connected components, $T_{e}^{+}\ni t(e)$ and $T_{e}^{-}\ni o(e)$. Set $T_{\geq e}:=T_{e}^{+}\sqcup\{e\}$ and $T_{\geq\bar{e}}:=T_{e}^{-}\sqcup\{\bar{e}\}$. A tree $T$ is uniquely geodesic, i.e., for all $v,w\in VT$ there is a unique reduced path $[v,w]$ from $v$ to $w$, which we call geodesic from $v$ to $w$. Recall that $[v,w]$ is the path of minimal length in $T$ from $v$ to $w$. Moreover, given $e,f\in ET$, there is a geodesic $(e_{1},\ldots,e_{n})$ in $T$ with $e_{1}=e$ and $e_{n}=f$ if, and only if, $f\in T_{\geq e}$. In general, for $t_{1},t_{2}\in T$ we denote by $[t_{1},t_{2}]$ the geodesic from $t_{1}$ to $t_{2}$ in $T$ (whenever it exists). Moreover, for non-empty subsets $X,Y\subseteq T$, denote by $\mathrm{Geod}_{T}(X\to Y)$ the set of all geodesics in $T$ from some $x\in X$ to some $y\in Y$. Finally, a ray in a tree is a sequence of edges $(e_{i})_{i\in\mathbb{Z}_{\geq 1}}$ such that $o(e_{i})\neq o(e_{j})$ and $t(e_{i})=o(e_{i+1})$, for all $i,j\in\mathbb{Z}_{\geq 1}$ with $i\neq j$.
@@
Let $\Gamma$ be a connected graph without $n$-cycles for every $n\geq 2$. Then $\Gamma$ has a unique maximal subtree: the subgraph $\Lambda$ obtained from $\Gamma$ by removing all its $1$-loops. In particular, for all $v,w\in V\Gamma$ the geodesic $[v,w]=(v=v_{0},e_{1},v_{1},\ldots,e_{n},v_{n}=w)$ in $\Lambda$ is the path of minimal length in $\Gamma$ from $v$ to $w$. Thus, if $v\neq w$ then $v_{i}\neq v_{j}$ for all $0\leq i,j\leq n$ with $i\neq j$.
@@ 3. Group actions on trees
In this section, we introduce the two main classes of group actions on trees considered in this paper. The first is the class of (P)-closed actions on trees (cf. Section 3.3). In Section 3.4, we briefly recall a local–to–global approach due C. Reid and S. Smith [17] to study these kinds of group actions, which will be largely used in the paper. We also add some new vocabulary (cf. Definition 3.4) that will be exploited in Section 4.3 to count geodesics on the tree of those actions. Section 3.5 introduces the second class we focus on, the one of weakly locally $\infty$-transitive actions on trees. Therein we discuss the connection with locally $\infty$-transitive actions (cf. Lemma 3.10), and provide a local characterisation and some explicit examples (cf. Proposition 3.11 and Example 3.12).
@@ 3.1. Group actions on trees
Let $T$ be a tree with $ET\neq\emptyset$ and without leaves (i.e., no vertices $v\in VT$ have $|o^{-1}(v)|=1$), and let $G$ be a topological group. A $G$-action $(G,T)$ on $T$ is a continuous group homomorphism $G\to\mathrm{Aut}(T)$ satisfying, for all $g\in G$ and $e\in ET$, that $g\cdot e\neq\bar{e}$ and, for all $v\in VT$, that $g\cdot v\neq v$ for some $g\in G$. In the literature, the latter two requirements are often added as separate conditions called “acting without edge inversions” and “acting without global fixed points”, respectively.
Given $t\in T$, denote by $G_{t}$ the stabiliser of $t$ in $G$. More generally, for every subset $X\subseteq T$, let $G_{X}$ denote the pointwise stabiliser of $X$. If $\mathfrak{p}=(e_{i})_{1\leq i\leq n}$ is a path, $G_{\mathfrak{p}}$ denotes the pointwise stabiliser of the set $\{e_{1},\ldots,e_{n}\}$.
An action $(G,T)$ is edge-transitive if $ET=G\cdot e\sqcup G\cdot\bar{e}$ for some (and hence every) $e\in ET$. Moreover, $(G,T)$ is locally $\infty$-transitive if, for every $v\in VT$ and $d\geq 0$, the stabiliser $G_{v}$ acts transitively on $\{\mathfrak{p}\in\mathrm{Geod}_{T}(v\to T)\mid\ell(\mathfrak{p})=d\}$ (cf. [5, §0.2]). One checks that locally $\infty$-transitive actions are edge-transitive. Examples of groups admitting a locally $\infty$-transitive action on a tree are the $k$-points of simple simply connected algebraic $k$-group of relative rank $1$, where $k$ is a non-Archimedean local field (cf. [20, pp. 91 and 95]), and the Burger–Mozes universal groups $U(F)$ associated to $2$-transitive groups $F\leq\mathrm{Sym}(\{1,\ldots,d\})$ acting on the barycentric subdivision of a $d$-regular tree (cf. the lines before [5, §3.1]).
@@ 3.2. The quotient graph and its standard edge weight
Let $(G,T)$ be a group action on a tree. The quotient graph $\Gamma=G\backslash T$ of the action is the graph with $V\Gamma:=G\backslash VT$, $E\Gamma:=G\backslash ET$ and, given $G\cdot e\in E\Gamma$, its origin is $G\cdot o(e)$, its terminus is $G\cdot t(e)$ and its inverse edge is $G\cdot\bar{e}$. One checks that these definitions are independent from the choice of $e$ in $G\cdot e$.
The assignment $\pi\colon t\in T\longmapsto G\cdot t\in\Gamma$ yields a graph epimorphism which is called the quotient map of $(G,T)$. The map $\pi$ entrywise extends to a map (denoted with the same symbol) from the set of all paths in $T$ to the set of all paths in $\Gamma$.
@@
Starting from a connected graph $\Gamma$ and a function $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}\cup\{\infty\}$, one always has a group action on a tree with quotient graph $\Gamma$ and standard edge weight $\omega$ (cf. Example 1.1).
@@
The study of the group $\mathrm{Aut}_{\pi}(T)$ as in Example 1.1 initiated the study of (P)-closed group actions on trees (cf. [2]), a class which stems from the more general class of group actions on trees with the Tits’ independence property (cf. [25, §4.2]). A group action on a tree $(G,T)$ is (P)-closed if $G\leq\mathrm{Aut}(T)$ is closed and, for every $e\in ET$,
@@
Note that the inclusion $\supseteq$ is automatic.
Proposition 3.1.
Let $(G,T)$ be a (P)-closed action on a tree and $(e_{1},\ldots,e_{n})$ be a geodesic in $T$ of length $n\geq 2$. Then, for every $k<n$ we have

### v2

@@ 2.1. Generalities
If there is no ambiguity, we denote a $1$-point set $\{x\}$ by the element $x$ itself. Given a set $X$, for every subset $A\subseteq X$ denote by $\mathbbm{1}_{A}$ the indicator function of $A$. For every set $X$, the symmetric group $\mathrm{Sym}(X)$ is always regarded as a topological group with the permutation topology, i.e., the topology generated by the local basis at $1$ given by all possible pointwise stabilisers of finite non-empty subsets of $X$.
@@ 2.2. Graphs
A graph (in the sense of J-P. Serre [27]) consists of a set $\Gamma=V\Gamma\sqcup E\Gamma$ partitioned into two subsets $V\Gamma$ and $E\Gamma$ (called the set of vertices and the set of edges of $\Gamma$, respectively), together with two maps $o,t\colon E\Gamma\to V\Gamma$ (called origin and terminus maps, respectively) and an involution $\overline{\cdot}\colon E\Gamma\to E\Gamma$ (called edge inversion) satisfying $\bar{e}\neq e$ and $o(\bar{e})=t(e)$, for every $e\in E\Gamma$. We introduce the following notation for a graph $\Gamma$.
@@
An orientation in a graph $\Gamma$ is a set $E\Gamma^{+}\subseteq E\Gamma$ satisfying $|\{e,\bar{e}\}\cap E\Gamma^{+}|=1$ for every $e\in E\Gamma$, and $E\Gamma=E\Gamma^{+}\cup\{\bar{e}\mid e\in E\Gamma^{+}\}$. A graph $\Gamma$ is non-empty if it has at least one vertex, and is locally finite if $|o^{-1}(c)|<\infty$ for all $c\in V\Gamma$.
A subgraph of a graph $\Gamma$ is a graph $\Lambda=V\Lambda\sqcup E\Lambda$ with $V\Lambda\subseteq V\Gamma$, $E\Lambda\subseteq E\Gamma$, and whose origin, terminus and inversion maps are obtained from the corresponding ones of $\Gamma$ by restriction. A subgraph $\Lambda$ of $\Gamma$ is proper if $\Lambda\neq\Gamma$.
A vertex $v$ in $\Gamma$ is said to be terminal (or a leaf) if $|o^{-1}(v)|=1$. An edge $e$ in $\Gamma$ with $o(e)=t(e)$ is called $1$-loop. A $n$-bouquet of loops (based at $c$) is a graph with one vertex $c$ and edge-set $\{a_{i},\bar{a}_{i}\mid 1\leq i\leq n\}$, where each $a_{i}$ is a $1$-loop starting at $c$. A $1$-segment is a graph $\Gamma$ with two distinct vertices and an edge-couple $\{e,\bar{e}\}$ connecting them.
@@
Let $\Gamma$ be a graph. A path in $\Gamma$ is a sequence of vertices and edges $\mathfrak{p}=(v_{0},e_{1},v_{1},\ldots,e_{n},v_{n})$, $n\geq 0$, with $o(e_{i})=v_{i-1}$ and $t(e_{i})=v_{i}$ for every $1\leq i\leq n$. Note that paths may contain repeated vertices and edges, and it might happen that $e_{i+1}=\bar{e}_{i}$ for some $i$. We say that $\mathfrak{p}$ starts at $v_{0}$ (or at $e_{1}$) and ends at $v_{n}$ (or at $e_{n}$), has reverse path $\overline{\mathfrak{p}}=(v_{n},\bar{e}_{n},v_{n-1},\ldots,\bar{e}_{1},v_{0})$, and length $n$ (written $\ell(\mathfrak{p})=n$). If $n\geq 1$, we may without ambiguity specify only the sequence of edges of $\mathfrak{p}$. If $n=0$, the $1$-term sequence $\mathcal{O}_{v_{0}}=(v_{0})$ is called the trivial path at $v_{0}$. Denote by $\mathcal{P}_{\Gamma}$ the set of all paths in $\Gamma$. Given non-empty subsets $X,Y\subseteq\Gamma$, let $\mathcal{P}_{\Gamma}(X\to Y)$ be the set of all paths in $\Gamma$ starting at some $x\in X$ and ending at some $y\in Y$. The product of two paths $\mathfrak{p}=(e_{1},\ldots,e_{m})$ and $\mathfrak{q}=(f_{1},\ldots,f_{n})$ is defined only if $t(e_{m})=o(f_{1})$ and it is the path $\mathfrak{p}\cdot\mathfrak{q}=(e_{1},\ldots,e_{m},f_{1},\ldots,f_{n})$. If $\mathfrak{p}$ is a path starting and ending at the same vertex, denote by $\mathfrak{p}^{d}$ the $d$-th power of $\mathfrak{p}$ with respect to the product defined before. A path $\mathfrak{p}$ is reduced if either $\ell(\mathfrak{p})=0$ or $\mathfrak{p}=(e_{1},\ldots,e_{n})$ and $e_{i+1}\neq\bar{e}_{i}$ for every $1\leq i\leq n-1$. For $n\geq 1$, an $n$-cycle is a reduced path $\mathfrak{p}=(e_{1},\ldots,e_{n})$ with $o(e_{1})=t(e_{n})$ and $t(e_{i})\neq t(e_{j})$ for all $1\leq i,j\leq n$ with $i\neq j$.
@@
A graph $\Gamma$ is connected if for all $v,w\in V\Gamma$ there is a path from $v$ to $w$. A subgraph of $\Gamma$ is a connected component if it is a maximal connected subgraph of $\Gamma$. A graph is the disjoint union of all its connected components. A tree $T$ is a connected graph with no $n$-cycles, for every $n\geq 1$. If $T$ is a tree and $e\in ET$, then the graph $T\setminus\{e,\bar{e}\}$ has two connected components, $T_{e}^{+}\ni t(e)$ and $T_{e}^{-}\ni o(e)$. Set $T_{\geq e}:=T_{e}^{+}\sqcup\{e\}$ and $T_{\geq\bar{e}}:=T_{e}^{-}\sqcup\{\bar{e}\}$. A tree $T$ is uniquely geodesic, i.e., for all $v,w\in VT$ there is a unique reduced path $[v,w]$ from $v$ to $w$, which we call geodesic from $v$ to $w$. Recall that $[v,w]$ is the path of minimal length in $T$ from $v$ to $w$. Moreover, given $e,f\in ET$, there is a geodesic $(e_{1},\ldots,e_{n})$ in $T$ with $e_{1}=e$ and $e_{n}=f$ if, and only if, $f\in T_{\geq e}$. In general, for $t_{1},t_{2}\in T$ we denote by $t_{1},t_{2}]$ the geodesic from $t_{1}$ to $t_{2}$ in $T$ (if it exists). Moreover, for non-empty subsets $X,Y\subseteq T$, denote by $\mathrm{Geod}_{T}(X\to Y)$ the set of all geodesics in $T$ from some $x\in X$ to some $y\in Y$.
A ray in a tree is a sequence of edges $(e_{i})_{i\in\mathbb{Z}_{\geq 1}}$ such that $o(e_{i})\neq o(e_{j})$ and $t(e_{i})=o(e_{i+1})$, for all $i,j\in\mathbb{Z}_{\geq 1}$ with $i\neq j$. Two rays $(e_{i})_{i\in\mathbb{Z}_{\geq 1}}$ and $(f_{i})_{i\in\mathbb{Z}_{\geq 1}}$ are equivalent if there is $j\in\mathbb{Z}_{\geq 1}$ such that $e_{i}=f_{i}$ for every $i\geq j$. Equivalence classes of rays in a tree are called ends.
@@
Let $\Gamma$ be a connected graph without $n$-cycles for every $n\geq 2$. Then $\Gamma$ has a unique maximal subtree: the subgraph $\Lambda$ obtained from $\Gamma$ by removing all its $1$-loops. In particular, for all $v,w\in V\Gamma$ the geodesic $v,w]=(v=v_{0},e_{1},v_{1},\ldots,e_{n},v_{n}=w)$ in $\Lambda$ is the path of minimal length in $\Gamma$ from $v$ to $w$. Thus, if $v\neq w$ then $v_{i}\neq v_{j}$ for all $0\leq i,j\leq n$ with $i\neq j$.
@@ 3. Group actions on trees
In this section, we introduce the two main classes of group actions on trees considered in this paper. The first is the class of (P)-closed actions on trees (cf. Section 3.3). In Section 3.4, we briefly recall a local–to–global approach due C. Reid and S. Smith [24] to study these kinds of group actions, which will be largely used in the paper. We also add some new vocabulary (cf. Definition 3.4) that will be exploited in Section 4.3 to count geodesics on the tree of those actions. Section 3.5 introduces the second class we focus on, the one of weakly locally $\infty$-transitive actions on trees. Therein we discuss the connection with locally $\infty$-transitive actions (cf. Lemma 3.10), providing a local characterisation and some explicit examples (cf. Proposition 3.11 and Examples 3.12, 3.13 and 3.14).
@@ 3.1. Group actions on trees
Let $G$ be a topological group and $T$ be a tree. A $G$-action $(G,T)$ on $T$ is a continuous group homomorphism $G\to\mathrm{Aut}(T)$ (recall that $\mathrm{Aut}(T)$ has the permutation topology, see Section 2.2). The action $(G,T)$ is said to be without edge inversions if $g\cdot e\neq\bar{e}$ for all $g\in G$ and $e\in ET$.
Convention 3.0.1.
In this paper, all actions on trees are assumed to be without edge inversions, and the relevant tree is assumed to have at least one edge.
Convention 3.0.1 allows us to have, as quotient of the action, a graph in the sense of J-P. Serre that has at least one edge. These (relatively mild) hypotheses are important for the introduction of certain linear operators associated the quotient graph (cf. Sections 5.4 and 5.5), whose role is key in our study of double-coset zeta functions.
Given $t\in T$, denote by $G_{t}$ the stabiliser of $t$ in $G$. More generally, for every subset $X\subseteq T$, let $G_{X}$ denote the pointwise stabiliser of $X$. If $\mathfrak{p}=(e_{i})_{1\leq i\leq n}$ is a path, $G_{\mathfrak{p}}$ denotes the pointwise stabiliser of the set $\{e_{1},\ldots,e_{n}\}$. If $\xi$ is an end of $T$ (regarded as a set of equivalent rays), we denote by $G_{\xi}$ its setwise stabiliser.
An action $(G,T)$ is arc-transitive if $ET=G\cdot e\sqcup G\cdot\bar{e}$ for some (and hence every) $e\in ET$. Moreover, $(G,T)$ is locally $\infty$-transitive if, for all $v\in VT$ and $d\geq 0$, the stabiliser $G_{v}$ acts transitively on $\{\mathfrak{p}\in\mathrm{Geod}_{T}(v\to T)\mid\ell(\mathfrak{p})=d\}$ (cf. [5, §0.2]). One checks that locally $\infty$-transitive actions are arc-transitive. Examples of groups admitting a locally $\infty$-transitive action on a tree satisfying Convention 3.0.1 are the $k$-points of simple simply connected algebraic $k$-groups of relative rank $1$, where $k$ is a non-Archimedean local field (cf. [27, pp. 91 and 95]), and the Burger–Mozes universal groups $U(F)$ associated to $2$-transitive groups $F\leq\mathrm{Sym}(\{1,\ldots,d\})$ acting on the barycentric subdivision of a $d$-regular tree (cf. the lines before [5, §3.1]).
@@ 3.2. The quotient graph and its standard edge weight
Let $(G,T)$ be a group action on a tree without edge inversions. The quotient graph $\T\Gamma=G\backslash$ T of the action is the graph with $\VTV\Gamma:=G\backslash V$T, $\ETE\Gamma:=G\backslash E$T and, given $G\cdot e\in E\Gamma$, its origin is $G\cdot o(e)$, its terminus is $G\cdot t(e)$ and its inverse edge is $G\cdot\bar{e}$. One checks that these definitions are independent of the choice of $e$ in $G\cdot e$.
The assignment $\pi\colon t\in T\longmapsto G\cdot t\in\Gamma$ yields a graph epimorphism which is called quotient map of $(G,T)$. The map $\pi$ entrywise extends to a map (denoted with the same symbol) from the set of all paths in $T$ to the set of all paths in $\Gamma$.
@@
Starting from a connected graph $\Gamma$ and a function $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}$, one always has a group action on a tree without edge inversions having quotient graph $\Gamma$ and standard edge weight $\omega$ (cf. Example 1.1).
@@
The study of the group $\mathrm{Aut}_{\pi}(T)$ as in Example 1.1 initiated the study of (P)-closed group actions on trees (cf. [2]), a class that stems from the more general class of group actions on trees with the Tits’ property (P) (cf. [32, §4.2]), which we now recall.
Let $(G,T)$ be a group action on a tree. Let $\ell=(v_{0},e_{1},v_{1},\ldots,e_{n},v_{n},\ldots)$, $1\leq n\leq+\infty$, be a geodesic or a ray in $T$ with vertices $(v_{n})_{n\geq 0}$ and edges $(e_{n})_{n\geq 1}$. Denote by $\pi_{\ell}\colon VT\to\{v_{n}\mid n\geq 0\}$ the closest point projection of the vertices of $T$ onto $\ell$. For every $n\geq 0$, the pointwise stabiliser $G_{\ell}$ induces a permutation group $(G_{\ell})_{n}\leq\mathrm{Sym}(\pi_{\ell}^{-1}(v_{n}))$. Therefore, one obtains a group monomorphism
$\Phi_{\ell}\colon G_{\ell}\longrightarrow\prod_{n\geq 0}(G_{\ell})_{n}.$
If $\Phi_{\ell}$ is surjective for every $\ell$, the action $(G,T)$ is said to have Tits’ property (P). Note that, if $\ell=(v_{0},e,v_{1})$ has only one edge $e$, the function $\pi_{\ell}$ maps every $v\in VT_{\geq\bar{e}}$ to $v_{0}=o(e)$ and every $v\in VT_{\geq e}$ to $v_{1}=t(e)$. Hence, it is easily checked that the surjectivity of $\Phi_{\ell}$ is equivalent to the requirement
@@
Following the characterisation of [24, Theorem 2.18], we say that the action $(G,T)$ is (P)-closed if it has Tits’ property (P) and $G\leq\mathrm{Aut}(T)$ is closed with respect to the permutation topology.
The following fact is well-known and is included for completeness.
Lemma 3.1.
Let $(G,T)$ be a (P)-closed action on a tree and $(e_{1},\ldots,e_{n})$ be a geodesic in $T$ of length $n\geq 2$. Then, for every $k<n$, we have